# FreqKV: Frequency Domain Key-Value Compression for Efficient Context Window Extension

### v1

@@ Abstract
Extending the context window in large language models (LLMs) is essential for applications involving long-form content generation. However, the linear increase in key-value (KV) cache memory requirements and the quadratic complexity of self-attention with respect to sequence length present significant challenges during fine-tuning and inference. Existing methods suffer from performance degradation when extending to longer contexts. In this work, we introduce a novel context extension method that optimizes both fine-tuning and inference efficiency. Our method exploits a key observation: in the frequency domain, the energy distribution of the KV cache is primarily concentrated in low-frequency components. By filtering out the high-frequency components, the KV cache can be effectively compressed with minimal information loss. Building on this insight, we propose an efficient compression technique, FreqKV, that iteratively compresses the increasing KV cache to a fixed size in the frequency domain, applicable to both fine-tuning and inference. FreqKV introduces no additional parameters or architectural modifications. With minimal fine-tuning, LLMs can learn to leverage the limited cache that is compressed in the frequency domain and extend the context window efficiently. Experiments on various long context language modeling and understanding tasks demonstrate the efficiency and efficacy of the proposed method.
@@ 1 Introduction
Large language models (LLMs) typically have a limited size of context window, which is pre-defined during the pre-training process. However, it is inevitable for LLMs to process sequences that exceed the preset context size. LLMs struggle to maintain their performance when generalized to longer contexts. Additionally, the computation cost of the self-attention mechanism [23] grows quadratically with the context length, meaning that doubling the context window results in a fourfold increase in the computational cost of attention modules.
For efficiency, existing efforts aim to compress the key-value (KV) cache for long contexts during inference. They evict [26, 17] or merge [29, 24] KV states of less important tokens following certain rules. They use attention scores to measure the importance and approximate the original full attention. However, while these methods provide an approximation of full computation on existing tokens through different strategies, they can not fully prevent performance degradation when decoding future tokens.
@@
Recent studies propose to fine-tune LLMs to longer contexts to extend the context window. LongLoRA [7] trains LLMs using shifted sparse attention. Despite training efficiency, their sparse attention fails to be applied during inference, and they still require the original attention on the full sequence. Concurrently, LoCoCo [5] and Activation Beacon [28] introduce additional modules to compress KV states. They incorporate the fine-tuned compressing pattern into the decoding procedure of LLMs.
@@
In the field of computer vision, studies have shown that low-frequency channels are more important for convolutional neural networks (CNNs) [27]. Moreover, Fourier Transformer [13] discards the high-frequency parts of the contexts and downsample the hidden states in the encoders. Inspired by these works, we seek to compress KV states in the frequency domain without the need for additional compression modules in decoder-only LLMs. We transform key states and value states in LLaMA-2-7b [22] from the time domain to the frequency domain for power spectrum analysis. As shown in Figure 1, the energy distribution increasingly concentrates on low-frequency components as the computation process progresses along the decoder layers. It suggests that high-frequency components, which occupy less energy and contribute less to the overall information, can be discarded without much performance loss, thereby enhancing computational efficiency.
@@
In this paper, we introduce FreqKV, an efficient context extension method that iteratively compresses key-value states in the frequency domain. Compression is triggered only when the KV cache reaches the predefined context window size. During each compression step, the low-frequency components of the KV states are preserved at a specified retaining ratio. Subsequent tokens are appended to the compressed cache until it is filled again. This ensures that the cached KV states that each query token can attend to is limited below the context window size. To reduce memory and computational costs, the compressed cache will be further compressed together with the incoming tokens. This iterative compression mechanism leads to an increased compression level of the earlier contexts as the sequence length grows. Without introducing additional compression modules, LLMs could learn to utilize the compressed cache efficiently when extending to longer contexts. FreqKV demonstrates comparable performance to other methods that employ full KV cache or additional compressors in long context language modeling. Furthermore, experiments on downstream tasks indicate that FreqKV surpasses recently studied KV compression methods in long-context understanding.
@@ 2 Related Work
To extend the context window of LLMs efficiently, researchers attempt to compress the KV cache as more tokens are fed into the model. One common approach is selective token eviction [26, 17, 6], where less significant tokens are discarded. Although the eviction strategies ensure that the size of KV cache involved in each decoding step does not exceed the pre-defined context window size, LLMs suffer from the permanent loss of the information associated with evicted tokens. To address this limitation, some researchers introduce cache merging techniques to approximate the original full attention of the existing contexts [29, 24, 25]. However, these inference methods often sacrifice performance for efficiency.
Context Extension for LLMs.
Recent advancements in context extension for LLMs have focused on efficiently scaling models to handle longer input sequences without significantly increasing computational costs. LongLoRA [7] employs shifted sparse attention during the parameter-efficient fine-tuning. However, this sparse attention mechanism is not applicable during inference, necessitating a return to the original full attention post-training. Other techniques, such as LoCoCo [5], integrate convolutional operations into LLMs for compressing long contexts. They fine-tune the compression modules together with LLMs. Landmark attention [19] uses landmark tokens to retrieve previous input blocks. Similarly, Activation Beacon [28] introduces a special token to represent the previous context for compression. However, they need a copy of multi-head attention, which amounts to approximately 2B for 7B models. In contrast, our proposed method achieves context extension without introducing any additional parameters.
@@
Learning in the frequency domain is a well-established technique to compress images and accelerate CNNs [11]. It has been observed that CNNs are more sensitive to low-frequency channels than high-frequency channels [27]. These works have inspired efforts to process natural language. FNet [16] enhances the efficiency of Transformer encoder architectures by replacing the self-attention layers with the Fourier transform to serve the purpose of mixing tokens. Additionally, Fourier Transformer [13] eliminates redundancies in the context through frequency domain processing within encoder architectures.
However, because of the auto-regressive nature, it remains unclear how to leverage frequency components for decoder-only Transformer, which is the main architecture of generative LLMs. To the best of our knowledge, FreqKV is the first work that explores compressing key-value states in the frequency domain for decoder-only LLMs.
@@ 3.1 Discrete Cosine Transform
The Discrete Cosine Transform (DCT) transforms a signal from the spatial domain (time or position) into the frequency domain. Several variants of the DCT exist, with DCT-II being the most common. For a real-value discrete signal ${\bm{X}}_{0:N-1}=[x_{0},\dots,x_{N-1}]$ of length $N$, it is defined as:
@@
The frequency components are expressed as a combination of the original signals. The values can be computed using the Fast Fourier Transform (FFT) with a complexity of $O(N\text{log}N)$. The amplitudes of frequency components are utilized in the power spectrum analysis to represent the energy or magnitude of components. The components of higher energy in the frequency domain indicate they are more informative [13].
@@ 3.2 Self-Attention
For the incoming token $x_{N}$, the prefilled $N$ tokens $X_{0:N-1}$ are utilized as the cache during decoding. Denote the hidden states of the $N+1$ tokens input to a specific layer of LLMs as ${\bm{H}}_{0:N}=[\displaystyle{\bm{h}}_{0},\cdots,\displaystyle{\bm{h}}_{N}]$. The query, key and value states of ${\bm{x}}_{N}$ are computed as follows::
${\bm{q}}_{N}=\displaystyle{\bm{h}}_{N}\displaystyle{\bm{W}}^{Q},\quad\displaystyle{\bm{k}}_{N}=\displaystyle{\bm{h}}_{N}\displaystyle{\bm{W}}^{K},\quad\displaystyle{\bm{v}}_{N}=\displaystyle{\bm{h}}_{N}\displaystyle{\bm{W}}^{V},$ (3)
where ${\bm{W}}^{Q},\displaystyle{\bm{W}}^{K},\displaystyle{\bm{W}}^{V}$ are the projection matrices for the query, key and value states, respectively. For simplicity, indices corresponding to layers and heads have been omitted.
The cached KV states for the previous $N$ tokens ${\bm{X}}_{0:N-1}$ are:
${\bm{K}}_{0:N-1}=\displaystyle{\bm{H}}_{0:N-1}\displaystyle{\bm{W}}^{K},\quad\displaystyle{\bm{V}}_{0:N-1}=\displaystyle{\bm{H}}_{0:N-1}\displaystyle{\bm{W}}^{V}.$ (4)
@@
where $d$ is the hidden dimension. $\oplus$ means the concatenation of the KV cache and KV states of $x_{N}$
@@ 4 FreqKV
Figure 1: The average power spectrums of key states and value states in different layers of LLaMA-2-7b. 1000 documents are sampled from CNN/Daily Mail [14]. We use DCT to transform key states and value states to the frequency domain, and average power spectrums over these samples and hidden dimensions.
@@ 4.1 Energy Concentration in the Frequency Domain
The figure shows that the energy of key states and value states is increasingly concentrated in the low-frequency components. The distributions of their power spectrums are similar since they are projected from the same hidden states as shown in Equation 3. Although there is no obvious energy concentration in the frequency domain for the initial embeddings of natural languages in the first layer, the model tends to aggregate energy in the low-frequency components along the decoding procedure. The observation of energy concentration suggests that we could maintain low-frequency components and filter out high-frequency components which could be redundant. Head-wise analysis of the power spectrum is provided in Appendix A.
@@ 4.2 KV Compression in the Frequency Domain
Figure 2: The illustration of our FreqKV. The KV cache will be compressed in an iterative manner as the cache reaches the context window size. Sink tokens remain uncompressed throughout the process. (a) The tokens after sink tokens will be compressed in the frequency domain and subsequent tokens will continue to get into the cache. (b) When the cache is filled again, the compressed tokens and incoming tokens will be compressed together. The compression is performed iteratively to extend the context window.
4.3 Context Extension with the Compressed KV
@@
For tokens within the context window, the standard attention will be conducted. For tokens out of the window, we will compress the cached KV states in the frequency domain. Because the retaining size $L$ is smaller than the cache size $N$, subsequent tokens could get in and fill the cache. We discard the original KV states of the prefilling tokens and maintain the compressed KV states. They will be concatenated with the KV states of the incoming tokens and compressed together when the cache is filled again. Tokens that appear earlier in the sequence undergo more iterations of compression as the context window expands, whereas less compression will be performed on the more recent tokens. With the iterative FreqKV, the KV cache size is not fixed and is reduced below the context window size during decoding. Since the compression is only performed when the sum of cached KV and incoming tokens reaches the preset size, the computation overhead of the compression could be negligible. For example, with the context size of 4096 and the retaining size of 2048, compression is performed every 2048 tokens for contexts exceeding the original window.
Recent work has found the phenomenon of attention sinks that LLMs tend to assign high attention scores to initial tokens [26, 12]. Therefore, we maintain these initial tokens uncompressed in the cache and only compress tokens that come after them.
During training and the prefilling stage of inference, the whole sentence is tokenized and fed into the model. The attention is computed chunk-wise interleaved with the compression operation. After each compression, $N-L-S$ incoming tokens are regarded as a chunk and fill the cache, with $S$ sink tokens uncompressed. As shown in Figure 2 (b), the incoming tokens in each chunk can not attend to the subsequent tokens. The newly incoming token $x_{M}$ will attend to $S$ sink tokens, $L$ compressed “tokens”, $M-N$ previous incoming tokens, and $x_{M}$ itself. The calculation of attention in Equation 9 can be reformulated as follows:
@@
It should be noted that key states in LLMs are often equipped with position embeddings like RoPE [21]. In FreqKV, key states are compressed and cached before applying RoPE. The cached key states are encoded with RoPE when conducting self-attention. They are allocated with position indices within the cache rather than the original sequence, which results in the context window extension without requiring position extrapolation or interpolation. Compressing key states with RoPE is explored in Appendix B.
@@ 5.1 Implementation
We conduct experiments on long context language modeling and understanding tasks with LLaMA-2-7b [22] and LLaMA-3-8b [1]. We maintain $S=4$ sink tokens uncompressed. The retaining ratio $\gamma$ in compression is set to 0.5. The preset context window size of LLaMA-2 is 4096, which is also the maximum KV cache size $N$. Therefore, the retaining size during each compression is $L=\gamma\cdot(N-S)=2046$. As long as the cache size reaches its capacity of 4096, the 4092 states since the $5$-th state in the cache will be compressed into 2046 states. The context size of LLaMA-3 is 8192. It is equipped with GQA (Grouped-Query Attention), which means it has a lower proportion of parameters for attention modules than LLaMA-2-Base/Chat-7b (Multi-Head Attention, MHA).
Minimal training is introduced to adapt models to this frequency-domain compression method. For long context language modeling, we fine-tune LLaMA-2-base on the RedPajama [8] pre-training dataset, extending the context window from 4K to 8K, 16K, and 32K. Perplexity (PPL) evaluation is conducted on PG-19 [20] and Proof-pile [2]. For long context understanding, the instruction following dataset LongAlpaca [7] is used for the supervised fine-tuning (SFT) of LLaMA-2-chat and LLaMA-3-instruct. The context window is extended from 4K to 8K, and from 8K to 16K respectively. Models are evaluated on LongBench [3] and Needle-in-a-Haystack [10]. Details of training settings are summarized in Appendix C.
Table 1: Perplexity evaluation on the test sets of PG-19 and Proof-pile. The superscript “*” means that we reproduce LoCoCo following their official code for evaluation. The results of full fine-tuning and LongLoRA are reported from [7].

### v2

@@ Abstract
Frequency-domain compression has proven effective in reducing redundancies for spatial signals. In this work, we propose FreqKV, a novel frequency domain key-value (KV) compression technique that enables efficient context window extension for decoder-only large language models (LLMs). Our approach is motivated by a key observation that, in the frequency domain, the energy distribution of the KV cache is predominantly concentrated in low-frequency components. By discarding high-frequency components, we achieve efficient compression of the KV cache with minimal information loss. FreqKV iteratively compresses the increasing KV cache to a fixed size in the frequency domain, allowing models to process lengthy contexts efficiently. Introducing no additional parameters or architectural modifications, FreqKV is applicable to both fine-tuning and inference. With minimal fine-tuning, LLMs can learn to leverage the limited cache that is compressed in the frequency domain and extend the context window. Experiments on a range of long context language modeling and understanding tasks demonstrate the efficiency and effectiveness of the proposed method.
@@ 1 Introduction
Large language models (LLMs) typically have a limited size of context window, which is pre-defined during the pre-training process. However, LLMs struggle to maintain their performance when processing sequences that exceed the preset context size. Additionally, the computational cost of the self-attention mechanism (Vaswani et al.,, 2023) grows quadratically with the context length, meaning that doubling the context window results in a fourfold increase in the computational cost of attention modules.
@@
Recent studies propose to fine-tune LLMs to longer contexts to extend the context window. LongLoRA (Chen et al.,, 2024) trains LLMs using shifted sparse attention. Despite training efficiency, their sparse attention fails to be applied during inference, and they still require the original attention on the full sequence. Concurrently, LoCoCo (Cai et al., 024a, ) and Activation Beacon (Zhang et al., 2024a, ) introduce additional modules to compress KV states and incorporate the fine-tuned compressing pattern into the decoding procedure of LLMs. However, the increasing memory requirement for additional parameters is inevitable.
For efficiency, existing efforts aim to compress the key-value (KV) cache for long contexts during inference. They evict (Xiao et al.,, 2024; Li et al.,, 2024; Jo et al.,, 2025) or merge (Zhang et al., 2024b, ; Wan et al.,, 2024) KV states of less important tokens following certain rules. They use attention scores to measure the importance and approximate the original full attention. However, while these methods provide an approximation of full computation on existing tokens, the information associated with the evicted tokens will be lost. They can not fully prevent performance degradation when decoding future tokens.
@@
In the field of computer vision, studies have shown that low-frequency channels are more important for convolutional neural networks (CNNs) (Xu et al.,, 2020). Moreover, Fourier Transformer (He et al.,, 2023) discards the high-frequency parts of the contexts and downsample the hidden states in the encoders. Inspired by these works, we seek to compress KV states in the frequency domain without the need for additional compression modules in decoder-only LLMs. We transform key states and value states in LLaMA-2-7b (Touvron et al.,, 2023) from the time domain to the frequency domain for power spectrum analysis. As shown in Figure 1, the energy distribution is increasingly concentrated on low-frequency components as the computation process progresses along the decoder layers. It suggests that high-frequency components, which occupy less energy and contribute less to the overall information, can be discarded without much performance loss, thereby enhancing computational efficiency.
@@
In this paper, we introduce FreqKV, an efficient context window extension method that iteratively compresses key-value states in the frequency domain. Compression is triggered only when the KV cache reaches the predefined context window size. During each compression step, KV states are converted to the frequency domain, where low-frequency components are retained before being transformed back to the time domain. Subsequent tokens are appended to the compressed cache until it is filled again. This ensures that the cached KV states that each query token can attend to are limited below the context window size, reducing memory and computational costs. The compressed cache undergoes further compression alongside incoming tokens. This iterative process progressively increases the compression ratio of earlier contexts as the sequence lengthens, while recent tokens remain less compressed, which is beneficial for autoregressive models to predicet next tokens. Without introducing additional compression modules, LLMs could learn to utilize the compressed cache efficiently when extending to longer contexts. FreqKV demonstrates comparable performance to other methods that employ full KV cache or additional compressors in long context language modeling. Furthermore, experiments on downstream tasks indicate that FreqKV surpasses recently studied KV compression methods in long context understanding.
@@ 2 Related Work
Context Extension for LLMs.
Recent advancements in context extension for LLMs have focused on efficiently scaling models to handle longer input sequences without significantly increasing computational costs. LongLoRA (Chen et al.,, 2024) employs shifted sparse attention during the parameter-efficient fine-tuning. However, this sparse attention mechanism is not applicable during inference, necessitating a return to the original full attention post-training. Other techniques, such as LoCoCo (Cai et al., 024a, ), integrate convolutional operations into LLMs for compressing long contexts. They fine-tune the compression modules together with LLMs. Landmark attention (Mohtashami and Jaggi,, 2023) uses landmark tokens to retrieve previous input blocks. Similarly, Activation Beacon (Zhang et al., 2024a, ) introduces a special token to represent the previous context for compression. However, they need a copy of multi-head attention, which amounts to approximately 2B for 7B models. In contrast, our proposed method achieves context extension without introducing any additional parameters.
@@
To extend the context window of LLMs efficiently, researchers attempt to compress the KV cache as more tokens are fed into the model. One common approach is selective token eviction (Xiao et al.,, 2024; Li et al.,, 2024; Cai. et al., 024b, ; Jo et al.,, 2025), where less significant tokens are discarded. Although the eviction strategies ensure that the size of KV cache involved in each decoding step does not exceed the pre-defined context window size, LLMs suffer from the permanent loss of the information associated with evicted tokens. To address this limitation, some researchers introduce cache merging techniques to approximate the original full attention of the existing contexts (Zhang et al., 2024b, ; Wan et al.,, 2024; Wang et al.,, 2024). However, these inference methods often sacrifice performance for efficiency.
@@
Learning in the frequency domain is a well-established technique to compress images and accelerate CNNs (Gueguen et al.,, 2018). It has been observed that CNNs are more sensitive to low-frequency channels than high-frequency channels (Xu et al.,, 2020). These works have inspired efforts to process natural language. FNet (Lee-Thorp et al.,, 2022) improves the efficiency of Transformer encoder architectures by replacing the self-attention layers with the Fourier transform to serve the purpose of mixing tokens. Additionally, Fourier Transformer (He et al.,, 2023) eliminates redundancies in the context through frequency domain processing within encoder architectures.
However, because of the autoregressive nature, it remains unclear how to leverage frequency components for decoder-only Transformer, which is the main architecture of generative LLMs. To the best of our knowledge, FreqKV is the first work that explores compressing key-value states in the frequency domain for decoder-only LLMs.
@@ 3.1 Discrete Cosine Transform
The Discrete Cosine Transform (DCT) transforms a signal from the spatial domain (time or position) into the frequency domain. Several variants of DCT exist, with DCT-II being the most common. For a real-value discrete signal ${\bm{X}}_{0:N-1}=[x_{0},\dots,x_{N-1}]$ of length $N$, it is defined as:
@@
The frequency components are expressed as a combination of the original signals. The values can be computed using the Fast Fourier Transform (FFT) with a complexity of $O(N\text{log}N)$. The amplitudes of frequency components are utilized in the power spectrum analysis to represent the energy or magnitude of components. The components of higher energy in the frequency domain indicate they are more informative (He et al.,, 2023).
@@ 3.2 Self-Attention
For the incoming token $x_{N}$, the prefilled $N$ tokens ${\bm{X}}_{0:N-1}$ are utilized as the cache during decoding. Denote the cached KV states for the previous $N$ tokens ${\bm{X}}_{0:N-1}$ as ${\bm{K}}_{0:N-1}$ and ${\bm{V}}_{0:N-1}$. For simplicity, indices corresponding to layers and heads have been omitted.
@@
where $d$ is the hidden dimension. ${\bm{q}}_{N},\displaystyle{\bm{k}}_{N}$ and ${\bm{v}}_{N}$ are the query, key and value states of $x_{N}$ respectively. $\oplus$ means the concatenation of the KV cache and KV states of $x_{N}$.
@@ 4 FreqKV
In this section, we present power spectrum analysis of KV states for LLMs. Building on this observation, FreqKV compresses KV states in the frequency domain and extends the context window via iterative compression.
@@
Figure 1: The average power spectrums of key states and value states in different layers of LLaMA-2-7b. 1000 documents are sampled from CNN/Daily Mail (Hermann et al.,, 2015). We use DCT to transform key states and value states to the frequency domain, and average power spectrums over these samples and hidden dimensions.
@@ 4.1 Energy Concentration in the Frequency Domain
The figure shows that the energy of key states and value states is increasingly concentrated in the low-frequency components. The distributions of their power spectrums are similar since they are projected from the same hidden states. Although there is no obvious energy concentration in the frequency domain for the initial embeddings of natural languages in the first layer, the model tends to aggregate energy in the low-frequency components along the decoding procedure. The observation of energy concentration suggests that we could maintain low-frequency components and filter out high-frequency components which could be redundant. Head-wise analysis of the power spectrum is provided in Appendix A.
@@ 4.2 KV Compression in the Frequency Domain
Figure 2: The illustration of our FreqKV. The KV cache will be compressed in an iterative manner as the cache reaches the context window size. Sink tokens remain uncompressed throughout the process. The tokens after sink tokens will be compressed in the frequency domain and subsequent tokens will continue to get into the cache. When the cache is filled again, the compressed tokens and incoming tokens will be compressed together. The compression is performed iteratively to extend the context window.
4.3 Context Extension via Iterative Compression
@@
For tokens within the context window, the standard attention will be conducted. For tokens out of the window, we will compress the cached KV states in the frequency domain. During each compression step, KV states are converted to the frequency domain via DCT, where low-frequency components are retained before applying IDCT to transform them back to the time domain. Because the retaining size $L$ is smaller than the cache size $N$, subsequent tokens could get in and fill the cache. The compressed cache will be concatenated with KV states of the incoming tokens and compressed together when the cache is filled again. Tokens that appear earlier in the sequence undergo more iterations of compression as the context window expands, whereas less compression will be performed on the more recent tokens. Since the compression is only performed when the sum of cached KV and incoming tokens reaches the preset size, the computation overhead of the compression could be negligible. For example, with the context size of 4096 and the retaining size of 2048, compression is performed every 2048 tokens for contexts exceeding the original window.
Recent work has found the phenomenon of attention sinks that LLMs tend to assign high attention scores to initial tokens (Xiao et al.,, 2024; Han et al.,, 2024). Therefore, we maintain these initial tokens uncompressed in the cache and only compress tokens that come after them.
During training and the prefilling stage of inference, the whole sentence is tokenized and fed into the model. The attention is computed chunk-wise interleaved with the compression operation. After each compression, $N-L-S$ incoming tokens are regarded as a chunk and fill the cache, with $S$ sink tokens uncompressed. As shown in Figure 2, the incoming tokens in each chunk can not attend to the subsequent tokens. The newly incoming token $x_{M}$ will attend to $S$ sink tokens, $L$ compressed “tokens”, $M-N$ previous incoming tokens, and $x_{M}$ itself. The calculation of attention in Equation 7 can be reformulated as follows:
@@
It should be noted that key states in LLMs are often equipped with position embeddings like RoPE (Su et al.,, 2023). In FreqKV, key states are compressed and cached before applying RoPE. The cached key states are encoded with RoPE when conducting self-attention. They are allocated with position indices within the cache rather than the original sequence, which results in the context window extension without requiring position extrapolation or interpolation. Compressing key states with RoPE is explored in Appendix B.
@@ 5.1 Implementation
We conduct experiments on long context language modeling and understanding tasks with LLaMA-2-7b (Touvron et al.,, 2023) and LLaMA-3-8b (AI@Meta,, 2024). We maintain $S=4$ sink tokens uncompressed. The retaining ratio $\gamma$ in compression is set to 0.5. The preset context window size of LLaMA-2 is 4096, which is also the maximum KV cache size $N$. Therefore, the retaining size during each compression is $L=\gamma\cdot(N-S)=2046$. As long as the cache size reaches its capacity of 4096, the 4092 states since the $5$-th state in the cache will be compressed into 2046 states. As for LLaMA-3, its context size is 8192. It is equipped with GQA (Grouped-Query Attention), which means it has a lower proportion of parameters for attention modules than LLaMA-2-Base/Chat-7b (Multi-Head Attention, MHA).
Minimal training is introduced to adapt models to this frequency-domain compression method. For long context language modeling, we fine-tune LLaMA-2-base on the pre-training dataset RedPajama (Computer,, 2023), extending the context window from 4K to 8K, 16K, and 32K. Perplexity (PPL) evaluation is conducted on PG-19 (Rae et al.,, 2019) and Proof-pile (Azerbayev et al.,, 2022). For long context understanding, the instruction following dataset LongAlpaca (Chen et al.,, 2024) is used for the supervised fine-tuning (SFT) of LLaMA-2-chat and LLaMA-3-instruct. The context window is extended from 4K to 8K, and from 8K to 16K respectively. Models are evaluated on LongBench (Bai et al.,, 2024) and Needle-in-a-Haystack (gkamradt,, 2023). Details of training settings are summarized in Appendix C.
Table 1: Perplexity evaluation on the test sets of PG-19 and Proof-pile. The superscript “*” means that we reproduce LoCoCo following their official code for evaluation. The results of full fine-tuning and LongLoRA are reported from (Chen et al.,, 2024).